OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (2, 4) and (4, 10)? y = −3x − 2 y = −3x + 2 y = 3x + 2 y = 3x − 2

OpenStudy (anonymous):

Y=3x-2

OpenStudy (anonymous):

why tho? @miroslava97

OpenStudy (anonymous):

Ok, use y2-y1/x2-x1, to get slope which is 3, then you choose a pair of points and plug it into the point slope form(y-y1=m(x-x1), y-4=3(x-2), which gives you the answer y=3x-2, best response?:)

OpenStudy (anonymous):

what bout the +4?